# HDRT: Infrared Capture for HDR Imaging

### v1

@@ 4. Experimental Results
DrTMO (Endo et al., 2017) 30.515 0.544 0.959 36.574 0.849 0.983 33.545 0.696 0.971
Deep Recursive HDRI (Lee et al., 2018) 32.997 0.728 0.959 41.139 0.938 0.988 37.068 0.833 0.974
HDRTVNet (Chen et al., 2021c) 31.074 0.633 0.953 39.739 0.802 0.984 35.407 0.717 0.969
LaNet (Yu et al., 2021b) 33.580 0.681 0.965 40.120 0.919 0.985 36.850 0.800 0.975
HDRCNN (Eilertsen et al., 2017) 35.515 0.703 0.966 42.737 0.872 0.988 39.126 0.788 0.977
Deep-HDR Reconstruction (Santos et al., 2020) 30.909 0.554 0.955 43.001 0.880 0.990 36.955 0.717 0.972
ICTCPNet (Huang et al., 2023) 21.371 0.253 0.938 31.271 0.629 0.974 26.321 0.441 0.956
ExpandNet (Marnerides et al., 2018) 25.633 0.415 0.938 39.288 0.882 0.985 32.461 0.649 0.962
HDRUNet (Chen et al., 2021a) 39.216 0.893 0.964 44.270 0.947 0.991 41.743 0.920 0.978
@@
4.1. Dataset
Existing RGBT datasets solely contain SDR images, typically focusing on the tasks of image fusion and object detection and segmentation. Existing HDR datasets, however, lack thermal infrared images. To study the effect of thermal imaging on HDR imaging performance, we have collected the first HDR and thermal dataset with aligned thermal and both SDR and HDR images. To investigate the performance under extreme visual conditions, our dataset comprises lighting conditions including underexposed and overexposed images.
For capturing IR images, we used an Optris PI 640i with 640$\times$480 pixel resolution and working wavelengths from 8 to 14 μm. This captures a temperature range from -20 °C to 100 °C. We employed a Canon EOS 5D Mark III for collecting the SDR images. Three different exposed SDR images are merged to recover HDR images using the method of Debevec et al. (Debevec and Malik, 1997). During the capturing process, the RGB and IR cameras are mounted on the same tripod to minimize the binocular camera registration error. After capturing, we manually register images from two cameras to ensure pixel-level alignment. Please refer to the supplementary for more statistics and details on the capture. We will make this dataset publicly available upon acceptance.
4.2. Comparison with Other Methods
We trained and tested the proposed method using PyTorch on an Intel Xeon W-2295 CPU and Nvidia Quadro 8000 GPU. We used Adam, with a learning rate of 4e-5, and 200 epochs. We used the MultiStepLR scheduler, and the learning rate was reduced by half every 20000 steps. The training set comprises of 8000 images; the validation set consists of 2000 images. To evaluate the performance compared to other methods, we utilized the perceptually uniform (pu) version of three commonly used metrics: peak signal-to-noise ratio (pu-PSNR), structural similarity index measure (pu-SSIM), and the visual saliency index (pu-VSI). These perceptually uniform metrics (Mantiuk and Azimi, 2021) are used as linear high dynamic range color values are non-linearly related to our perception of visible differences, so we converted absolute high dynamic range linear color values into perceptually uniform values.
As our approach is designed for single time capture (i.e. visible and IR information is captured simultaneously), we compared with the state-of-the-art single image HDR reconstruction methods and viewed multi frame capture approaches as being out of scope (e.g. multiple exposure video reconstruction approaches or flash/no flash pairs). We therefore compared with the following methods: DrTMO (Endo et al., 2017), Deep Recursive HDRI (Lee et al., 2018), HDRTVNet (Chen et al., 2021c), LaNet (Yu et al., 2021b), HDRCNN (Eilertsen et al., 2017), Deep-HDR Reconstruction (Santos et al., 2020), ICTCPNet (Huang et al., 2023), ExpandNet (Marnerides et al., 2018), and HDRUNet (Chen et al., 2021a). Table 1 presents the results of our HDRT dataset. Our analysis divided the table into three groups of images where HDR imaging is required, over-exposed images, under-exposed images, and a combination of these images. The proposed method obtains the best performance in all three groups.
Figure 6. Examples of HDRTNet with and without infrared images.
Table 2. Ablation study of the proposed method on our HDRT dataset. “√” indicates the technique is adopted, “×” indicates not, and “-” means not applicable.
@@ 4.2. Comparison with Other Methods
4.3. Ablation Study
To demonstrate the effectiveness of our major modules, we implemented three variations: baseline RGB without IR, pixel-level rather than feature-level fusion, and combined rather than separate training of the branches. Baseline RGB without IR refers to only using a standard U-Net (Ronneberger et al., 2015) to recover HDR images from SDR images (referred to as “RGB”). Pixel-level fusion also uses a standard U-Net, but different from RGB this method adopts IR image data as the fourth channel of the input, from which the network can extract information from the IR domain (referred to as “Pixel”). Combined training refers to fusing the features from IR and SDR images in feature space, but training together with the HDR branch rather than separately training and reconstructing RGB images (referred to as “Combined (Naive)”).
The comparison results on the HDRT dataset are shown in Table 2. Since the feature-level fusion and IR branch cannot operate with only RGB, we excluded these methods in Table 2.
RGB Baseline. Comparing “Pixel” with the baseline “RGB”, Pixel exceeds RGB in pu-PSNR, pu-SSIM, and pu-VSI by 0.415, 0.011, and 0.005, respectively. This is because, as discussed before, due to sensor limitations SDR images lose information compared with HDR images. The inclusion of data from the IR domain adds more details, as is illustrated in Figure 6.
Pixel-Level Fusion. However, comparing HDRTNet with “Pixel”, HDRTNet still exceeds Pixel in pu-PSNR, pu-SSIM, and pu-VSI by 0.530, 0.005, and 0.003, respectively. This is because, without a separate sub-network handling IR image data, the pixel-level fusion is prone to ignoring heterogeneous regions in IR images, resulting in uncompleted info from IR domain being extracted. As shown in Figure 7 (a), Pixel is visually better than RGB but still has some artifacts compared to HDRTNet.
Another problem with pixel-level fusion is that it is not robust enough for binocular camera registration errors. As shown in Figure 7 (b), IR and RGB images are not completely registered, and a clear deviation in their positions can be seen in the second column of fused images. The results of Pixel clearly show this ghosting, but HDRTNet does not. As a result, pixel-level fusion produces more blurry images due to tiny registration errors.
@@ 4.3. Ablation Study
Figure 7. Visual comparison between the pixel-level fusion and our HDRTNet.
Naive Combined Training. Comparing “Combined (Naive)” with “Pixel”, Combined exceeds Pixel in pu-PSNR and pu-SSIM by 0.134 and 0.001, respectively, which means that the feature-level fusion can help resolve the problem of output instability. As seen in Figure 8 (a), the output of Combined is very smooth and does not produce artifacts like Pixel. However, HDRTNet still exceeds Combined in pu-PSNR, pu-SSIM, and pu-VSI by 0.396, 0.005, and 0.002, respectively. This is because visually irrelevant features reduce the resulting quality of the HDR image, as shown in Figure 8 (b). In Figure 8 (b), the output from Combined is less saturated compared to the ground truth, while HDRTNet more closely resembles the color information present. Compared with the ground truth, the image generated by Combined is more like a fusion of the RGB and IR images (the first column of Figure 8 (b)), in which the colors are whiter and simpler than ground truth images. However, the IR branch in HDRTNet is trained to reconstruct RGB images, which can avoid extracting visual-irrelevant features. Therefore, HDRTNet is not easily affected by visually irrelevant features in the IR domain.
(a) Visual comparison between Pixel, Combined, and our HDRTNet.
@@
Figure 8. Visual comparison between the feature-level fusion and HDRTNet.
4.4. Visual Performance
@@
We also demonstrate qualitative results by comparing the proposed HDRTNet with the state-of-the-art and the original HDR frames. The input for all methods are the SDR images. All results are grouped by images containing mostly over-exposed and under-exposed regions. All HDR images use the Durand-Evrard tone mapping operator for a fair visual comparison.
@@
The first five columns in Figure 9 show the over-exposure cases. The main challenge in these scenarios is to recover cloud details in the first three columns and to restore missing details around the lights in the fourth and fifth columns. With additional information from the IR data, HDRTNet produces images closest to the ground truth image. In Figure 9, the last four columns show under-exposure cases, which makes it challenging to restore the building details in the sixth and seventh columns and the dark details of the nighttime images in the eighth and last columns. Again, HDRTNet utilizes IR information to generate HDR images closes to the ground truth images. For more visual results, please see the supplementary.
4.5. Larger-resolution Images
To demonstrate the effectiveness of our adaptations for larger-resolution images, we compared using our original model directly to process larger-resolution images with the adapted model, and the results are shown in Table 3. Here, the baseline is our original method using a resolution of 256$\times$256, which achieves the best performance in three metrics. Compared to the second column of Table 3 with the first column, the performance of the original method at 512$\times$512 reduces by 1.659, 0.016, and 0.003 in pu-PSNR, pu-SSIM, and pu-VSI, respectively.
This is because, when directly using U-Net to process larger-resolution images, the receptive field does not match. As shown in Figure 11 (a), larger-resolution images output by U-Net have artifacts, especially in red circles. Compared to the last column of Table 3 with the first column, the performance of the adapted method slightly reduces by 0.113 and 0.001 in pu-PSNR and pu-SSIM. In Figure 11 (a), larger-resolution images output by our adapted model do not have these artifacts, which demonstrates the effectiveness of the pixel-unshuffle handling larger-resolution images.
Compared to the last column of Table 3 with the third column, the performance of the adapted method increases by 0.091 and 0.004 in pu-PSNR and pu-SSIM, respectively. This is because, without the smooth gradient loss, the image output by the adapted model has grid patterns due to the surrounding four points being generated separately, as shown in Figure 11 (b). However, with the smooth gradient loss, the model can reduce the gradient generated in the horizontal and vertical coordinates, promoting image smoothness, as shown in Figure 11 (b).
Table 3. Comparison of the large-resolution images.
Methods pu-PSNR pu-SSIM pu-VSI
the original HDRTNet under 256$\times$256 46.033 0.954 0.994
the original HDRTNet under 512$\times$512 44.374 0.938 0.991
the adapted HDRTNet w/o the smooth gradient loss under 512$\times$512 45.829 0.949 0.994
the adapted HDRTNet w/ the smooth gradient loss under 512$\times$512 45.920 0.953 0.994
4.6. Discussion
@@
Our proposed method performs well in enhancing HDR in over- or under-exposed conditions. However, we still need to consider how to utilize IR info to improve HDR when the lighting conditions are completely well-exposed. Furthermore, some materials block IR transmission while letting through visible light. Our method is likely to be less effective in these cases. We currently use two cameras separately to capture IR and SDR images. Binocular errors may occur when there are far and near objects in the scenario, as shown in Figure 7 (b). In such cases, precise registration of IR and RGB images is necessary. Alternatively, a new registration pre-processing model can be developed to ensure accurate registration before feeding two images into the network, or ultimately, the IR and RGB sensors can be integrated into one camera, which requires hardware improvements.
5. Conclusion
In this paper, we have proposed, HDRT, a novel approach for HDR imaging via the use of thermal imaging. Our approach fills in the missing information in singly-exposed SDR images via leveraging information present in the thermal image. This combines features from both the IR and RGB data in a U-Net to create a resulting HDR image.
@@ 5. Conclusion
Extensive experiments have shown the effectiveness of our proposed method which outperforms other state-of-the-art methods and, in particular, excels at well under- and over-exposed images, in which there is information loss due to adverse lighting environments and sensor limitations. In addition, ablation experiments have demonstrated that our novel modules for feature-level fusion and infrared feature extraction can effectively extract infrared features and avoid HDR performance degradation caused by modality differences between infrared and visible light.
To summarize, our work proposes and validates the use of thermal infrared imaging to capture HDR images, and we hope future work will build on our pipeline for HDR content capturing.
Figure 9. A visual comparison of all tested methods.
Figure 10. Another visual comparison of all tested methods.
(a) A visual comparison of the original HDRTNet with resolutions of 256$\times$256 and 512$\times$512 and the adapted HDRTNet with a resolution of 512$\times$512.
(b) Examples of the adapted model with and without the smooth gradient loss.
Figure 11. Visual results of the adapted HDRTNet for handling large-size images.

### v2

@@ 4.1 Methodology
DrTMO [43] 30.515 0.544 0.959 36.574 0.849 0.983 33.545 0.696 0.971
Deep Recursive HDRI [27] 32.997 0.728 0.959 41.139 0.938 0.988 37.068 0.833 0.974
HDRTVNet [44] 31.074 0.633 0.953 39.739 0.802 0.984 35.407 0.717 0.969
LaNet [45] 33.580 0.681 0.965 40.120 0.919 0.985 36.850 0.800 0.975
HDRCNN [6] 35.515 0.703 0.966 42.737 0.872 0.988 39.126 0.788 0.977
Deep-HDR Reconstruction [46] 30.909 0.554 0.955 43.001 0.880 0.990 36.955 0.717 0.972
ICTCPNet [47] 21.371 0.253 0.938 31.271 0.629 0.974 26.321 0.441 0.956
ExpandNet [41] 25.633 0.415 0.938 39.288 0.882 0.985 32.461 0.649 0.962
HDRUNet [48] 39.216 0.893 0.964 44.270 0.947 0.991 41.743 0.920 0.978
@@
4.2 Comparison with Other Methods
We trained and tested the proposed method using PyTorch on an Intel Xeon W-2295 CPU and Nvidia Quadro 8000 GPU. We used Adam, with a learning rate of 4e-5, and 200 epochs. We used the MultiStepLR scheduler, and the learning rate was reduced by half every 20000 steps. The training set comprises of 8000 images; the validation set consists of 2000 images. To evaluate the performance compared to other methods, we utilized the perceptually uniform (pu) version of three commonly used metrics: peak signal-to-noise ratio (pu-PSNR), structural similarity index measure (pu-SSIM), and the visual saliency index (pu-VSI). These perceptually uniform metrics [49] are used as linear high dynamic range color values are non-linearly related to our perception of visible differences, so we converted absolute high dynamic range linear color values into perceptually uniform values.
As our approach is designed for single time capture (i.e. visible and IR information is captured simultaneously), we compared with the state-of-the-art single image HDR reconstruction methods and viewed multi frame capture approaches as being out of scope (e.g. multiple exposure video reconstruction approaches or flash/no flash pairs). We therefore compared with the following methods: DrTMO [43], Deep Recursive HDRI [27], HDRTVNet [44], LaNet [45], HDRCNN [6], Deep-HDR Reconstruction [46], ICTCPNet [47], ExpandNet [41], and HDRUNet [48]. Tab. 3 presents the results of our HDRT dataset. Our analysis divided the table into three groups of images where HDR imaging is required, over-exposed images, under-exposed images, and a combination of these images. The proposed method obtains the best performance in all three groups.
Figure 7: Examples of HDRTNet with and without infrared images.
Table 4: Ablation study of the proposed method on our HDRT dataset. “√” indicates the technique is adopted, “×” indicates not, and “-” means not applicable.
@@ 4.2 Comparison with Other Methods
4.3 Ablation Study
To demonstrate the effectiveness of our major modules, we implemented three variations: baseline RGB without IR, pixel-level rather than feature-level fusion, and combined rather than separate training of the branches. Baseline RGB without IR refers to only using a standard U-Net [50] to recover HDR images from SDR images (referred to as “RGB”). Pixel-level fusion also uses a standard U-Net, but different from RGB this method adopts IR image data as the fourth channel of the input, from which the network can extract information from the IR domain (referred to as “Pixel”). Combined training refers to fusing the features from IR and SDR images in feature space, but training together with the HDR branch rather than separately training and reconstructing RGB images (referred to as “Combined (Naive)”).
The comparison results on the HDRT dataset are shown in Tab. 4. Since the feature-level fusion and IR branch cannot operate with only RGB, we excluded these methods.
RGB Baseline. Comparing “Pixel” with the baseline “RGB”, Pixel exceeds RGB in pu-PSNR, pu-SSIM, and pu-VSI by 0.415, 0.011, and 0.005, respectively. This is because, as discussed before, due to sensor limitations SDR images lose information compared with HDR images. The inclusion of data from the IR domain adds more details, as is illustrated in Fig. 7.
@@ 4.3 Ablation Study
Figure 8: Visual comparison between the pixel-level fusion and our HDRTNet.
Pixel-Level Fusion. However, comparing HDRTNet with “Pixel”, HDRTNet still exceeds Pixel in pu-PSNR, pu-SSIM, and pu-VSI by 0.530, 0.005, and 0.003, respectively. This is because, without a separate sub-network handling IR image data, the pixel-level fusion is prone to ignoring heterogeneous regions in IR images, resulting in uncompleted info from IR domain being extracted. As shown in Fig. 8 (a), Pixel is visually better than RGB but still has some artifacts compared to HDRTNet.
Another problem with pixel-level fusion is that it is not robust enough for binocular camera registration errors. As shown in Fig. 8 (b), IR and RGB images are not completely registered, and a clear deviation in their positions can be seen in the second column of fused images. The results of Pixel clearly show this ghosting, but HDRTNet does not. As a result, pixel-level fusion produces more blurry images due to tiny registration errors.
Naive Combined Training. Comparing “Combined (Naive)” with “Pixel”, Combined exceeds Pixel in pu-PSNR and pu-SSIM by 0.134 and 0.001, respectively, which means that the feature-level fusion can help resolve the problem of output instability. As seen in Fig. 9 (a), the output of Combined is very smooth and does not produce artifacts like Pixel. However, HDRTNet still exceeds Combined in pu-PSNR, pu-SSIM, and pu-VSI by 0.396, 0.005, and 0.002, respectively. This is because visually irrelevant features reduce the resulting quality of the HDR image, as shown in Fig. 9 (b). In Fig. 9 (b), the output from Combined is less saturated compared to the ground truth, while HDRTNet more closely resembles the color information present. Compared with the ground truth, the image generated by Combined is more like a fusion of the RGB and IR images (the first column of Fig. 9 (b)), in which the colors are whiter and simpler than ground truth images. However, the IR branch in HDRTNet is trained to reconstruct RGB images, which can avoid extracting visual-irrelevant features. Therefore, HDRTNet is not easily affected by visually irrelevant features in the IR domain.
(a) Visual comparison between Pixel, Combined, and HDRTNet.
@@
Figure 9: Visual comparison between the feature-level fusion and HDRTNet.
Loss Weight. In order to test the impact of different weights $\alpha$ and $\beta$ on HDR performance, we use HDRTNet with different weights to train and reconstruct HDR images; the results are shown in Tab. 5. The results show that the performance is best when $\alpha$ is set to 1.0 and $\beta$ is set to $1\textrm{e}-5$.
Table 5: Ablation study on $\alpha$ and $\beta$.
$\alpha$ $\beta$ pu-PSNR pu-SSIM pu-VSI
e-1 e-6 44.292 0.933 0.991
e-1 e-5 44.516 0.942 0.991
e-1 e-4 44.292 0.933 0.991
1 e-6 44.802 0.940 0.991
1 e-5 45.359 0.949 0.994
1 e-4 44.811 0.940 0.991
e+1 e-6 44.506 0.936 0.991
e+1 e-5 44.435 0.934 0.991
e+1 e-4 44.895 0.943 0.991
4.4 Visual Performance
@@
This section demonstrates qualitative results by comparing the proposed HDRTNet with the state-of-the-art and the original HDR frames. The input for all methods are the SDR images. All results are grouped by images containing mostly over-exposed and under-exposed regions. The results can be seen in Fig. 10. All HDR images use the Durand tone mapping operator [51] for a fair visual comparison.
@@
The first five columns in Fig. 10 show the over-exposure cases. The main challenge in these scenarios is to recover cloud details in the first three columns and to restore missing details around the lights in the fourth and fifth columns. With additional information from the IR data, HDRTNet produces images closest to the ground truth image.
In Fig. 10, the last four columns show under-exposure cases, which makes it challenging to restore the building details in the sixth and seventh columns and the dark details of the nighttime images in the eighth and last columns. Again, HDRTNet utilizes IR information to generate HDR images closes to the ground truth images.
We also present a series of further results showcasing HDR images under various exposure settings. Included are comparisons between HDR images recovered by the proposed HDRTNet, the ground-truth HDR images, and counterparts generated by other state-of-the-art ITMOs. The results can be seen in Fig. 11. Our approach is closest to ground truth across all exposure values.
Furthermore, the effectiveness of the proposed HDRTNet can be extended to a diverse array of scenes, highlighting the robust generalization capability across different scenarios, as is shownin Fig. 12. These visual results demonstrate the strong HDR imaging performance of our HDRTNet and its broad applicability.
Figure 10: A visual comparison of all tested methods (tone-mapped).
Figure 11: A visual comparison of all tested methods (exposure blankets).
Figure 12: Visual results of the proposed HDRTNet in different scenarios.
4.5 Discussion
@@
Our proposed method performs well in enhancing HDR in over- or under-exposed conditions. However, we still need to consider how to utilize IR info to improve HDR when the lighting conditions are completely well-exposed. Furthermore, some materials block IR transmission while letting through visible light. Our method is likely to be less effective in these cases. We currently use two cameras separately to capture IR and SDR images. Binocular errors may occur when there are far and near objects in the scenario, as shown in Fig. 8 (b). In such cases, precise registration of IR and RGB images is necessary. Alternatively, a new registration pre-processing model can be developed to ensure accurate registration before feeding two images into the network, or ultimately, the IR and RGB sensors can be integrated into one camera, which requires hardware improvements.
5 Conclusion
In this paper, we have introduced HDRT, a novel dataset for imaging which consists of 10K paired thermal infrared and visible light images. Our dataset provides a comprehensive collection of images captured under diverse lighting conditions, which serves as a valuable resource for advancing HDR and IR imaging research.
@@ 5 Conclusion
Extensive experiments utilizing the HDRT dataset have demonstrated its effectiveness in improving HDR imaging techniques. Our proposed method which outperforms other state-of-the-art methods and, in particular, excels at well under- and over-exposed images, in which there is information loss due to adverse lighting environments and sensor limitations. In addition, ablation experiments have demonstrated that our novel modules for feature-level fusion and infrared feature extraction can effectively extract infrared features and avoid HDR performance degradation caused by modality differences between infrared and visible light.
The HDRT dataset not only provides a solid foundation for current IR-guided HDR imaging research but also holds significant potential for a wide range of applications, including other HDR methods, multi-modal fusion, domain adaptation, and advanced image reconstruction tasks. To summarize, our work proposes and validates the use of thermal infrared imaging to capture HDR images, and we hope future work will build on our pipeline for HDR content capturing.